AAPL wrote:A set of pictures of butterflies contains 10 pictures. Jim bought 3 of the pictures. If 2 pictures are to be picked out from the 10 pictures, what is the probability that both pictures are no those that were already bought by Jim?
A. 14/15
B. 3/5
C. 6/13
D. 7/15
E. 7/10
P(1st picture selected has not already been purchased) = 7/10. (Of the 10 pictures, 3 have already been purchased, so 7 have not.)
P(2nd picture selected has not already been purchased) = 6/9. (Of the 9 remaining pictures, 3 have already been purchased, so 6 have not.)
To combine these probabilities, we multiply:
7/10 * 6/9 = 7/15.
The correct answer is
D.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.
As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.
For more information, please email me (Mitch Hunt) at
[email protected].
Student Review #1
Student Review #2
Student Review #3